# Experimental evidence for the physical delocalization of individual photons in an interferometer

### v1

@@ 5 Discussion
Our results show how individual photons detected in the outputs of an interferometer are physically delocalized as they propagate through the interferometer, where the concept of delocalization is explained in terms of the empirical evidence provided by local interactions within the paths. By ensuring that the polarization flip rate is the same for all photons localized in only one of the paths, we identify an experimentally observable phenomenon directly associated with the delocalization of the particles between the two paths. As shown in Fig. 5, this phenomenon is consistent with the fact that particles are localized in one of the paths whenever they are detected in that paths. However, the situation changes when the particles are detected in the output ports of the interferometer instead. The majority of photons detected in the output port favored by constructive interference experiences a proportional combination of both local effects, indicating a corresponding delocalization of the photons within the interferometer. The stronger the interference effect, the more equal the distribution of each photon between the paths. This effect is consistent with the expectation that interference works best when the intensities of both beams are equal. Photons detected in constructive interference thus behave very much like classical wave intensities. However, their detection in the output port is still a particle-like all-or-nothing event, as demanded by the statistical interpretation of quantum mechanics. It is therefore necessary to also consider the much rarer case of photons detected in the port in which destructive interference reduces the detection probability. Since the average rate of polarization flips over all photons cannot be changed by the final measurement, the reduction of the flip rate by delocalization must be associated with an enhancement of the flip rate by the opposite effect of super-localization. The experimental results shown in Fig. 6 confirm this phenomenon. Whenever a photon is detected in destructive interference, it experiences a polarization rotation that is much larger than the local rotations. This indicates that the effects of the opposite rotations add up, suggesting a reversal of sign caused by a negative contribution in one path, and a corresponding increase in the contribution from the other path. In terms of classical wave propagation, one might be tempted to interpret this as an amplitude fluctuation that changed the sign of the amplitude. However, it should be kept in mind that such an overly classical interpretation could not be reconciled with the detection of individual particles. Instead, we should focus on the observation that detection in the unlikely output port is a extreme fluctuation and can therefore be correlated with equally extreme fluctuations of delocalization.
@@
To summarize our results, we have shown that the physical delocalization of individual particles inside an interferometer can be determined experimentally by using sufficiently weak interactions. The results show that particles are only localized when the final measurement is not sensitive to interference effects. These findings have serious implications for the way in which we think about quantum superpositions. In the absence of empirical evidence to the contrary, it is indeed tempting to claim that each component of a quantum state represents the outcome of a measurement, whether the measurement is performed or not. It is therefore important to recognize that our results demonstrate the fallacy of this counterfactual assumption. In the formalism of quantum mechanics, the role of state vectors and their superpositions is ambiguous. Only an actual experiment can identify the physical meaning of such concepts. Misunderstandings may have arisen from a rather hasty identification of physical properties with the eigenstates of operators that represent them in the formalism. As our previous work has shown[38, 39, 40], the experiment presented here is consistent with the standard formalism of quantum mechanics, and the definition of delocalization used here can be formalized in terms of the appropriate operator ordering (details in the appendix C). In the present work, we have highlighted the experimental realization because a purely mathematical definition of fundamental concepts would seem arbitrary and carry very little practical meaning. The problem of the conventional approach to quantum mechanics is that its abstractions hide the wide range of possibilities by which we can explore nature. The present work solves this problem for the case of individual particles in an interferometer, paving the way towards a better practical understanding of all quantum phenomena.
@@
Acknowledgements H.F.H. was supported by ERATO, Japan Science and Technology Agency (JPMJER2402)
Appendix A Reduction of systematic effects
A.1 Elliptical polarization components
@@
We found that imperfections of the BS used in the experiment introduced a small amount of elliptical polarization, corresponding to an additional phase shift $\phi_{HV}$ between the H and the V components. If this phase shift is the same for both paths, it causes a systematic error in the $P(H|\pm)$ measurements. To cancel out the elliptical components in the transmitted and reflected output paths, two HWPs placed in the interferometer were individually tilted towards the optical axis. The magnitude of the elliptical component $\langle S_{RL}\rangle$ was directly measured at the $+$ and the -- ports for path 1 and path 2 by blocking the respective other path in the interferometer. The adjustment of each HWP and the measurement of the elliptical component were iterated to reduce the sum value of the two elliptical polarization components. The measurement results for $\langle S_{RL}\rangle$ obtained with the configuration used in the experiment are summarized in Supplementary Table 1.
Supplementary Table 1: Summary of measurement results
@@
$+$ port $0.0759\pm 0.0007$ $-0.0604\pm 0.0006$ $0.0155\pm 0.0009$ $3.52\pm 0.2\degree$
-- port $0.0695\pm 0.0007$ $-0.0698\pm 0.0008$ $-0.0003\pm 0.0011$ $-0.07\pm 0.2\degree$
@@
The remaining sum value of the elliptical components $\Delta\langle S_{RL}\rangle$ of the transmitted and the reflected paths results in a phase offset in the horizontal axis of Fig. 5. The phase offset $\Delta\phi$ in Supplementary Table 1 was evaluated using the relation $\Delta\langle S_{RL}\rangle/2\theta_{0}$, where the elliptical component can be expressed as $\sin 2\theta_{0}\sin\phi_{HV}\approx 2\theta_{0}\phi_{HV}$. The overall effect of the systematic error is sufficiently small, but it is noticeable that the outcomes of the $P(H|-)$ measurement have a lower systematic error than the outcomes of the $P(H|+)$ measurement.
@@
The difference between the systematic errors is particularly noticeable in Fig. 5, where the phase offset $\Delta\phi$ shifts the crossing points of the single-path measurements and the interference measurements. For $P(H|-)$, the crossing point is almost precisely at $\phi=90\degree$, whereas for $P(H|+)$ it is not. The mismatch for $P(H|+)$ shown in Fig. 5 (b) is consistent with the offset angle of $\Delta\phi=3.52\degree$, which roughly corresponds to 2/3 of one step of the data points shown in Fig. 5 (b).
A.2 Additional polarization rotation by the final BS
@@
Another systematic effect originates from an additional polarization rotation caused by an unintended polarization dependence of transmission and reflection at the BS. This additional polarization rotation introduces a difference in the magnitude of the rotation observed in the two output ports that is not related to the delocalization of the photon. To avoid this systematic effect, the orientations of the GTs in the output ports were slightly adjusted so that the values of $P(H|1)$ and $P(H|2)$ observed when the other path was blocked had the same values in both output ports. This adjustment is necessary to satisfy the condition that the magnitude of the polarization rotation is determined by the opposite polarization rotations in the paths only, with no systematic dependence on the output path. To evaluate the compensation angle, we measured the dependence of $P(H|1)$ and $P(H|2)$ on the angle of the GT $\theta_{GT}$ by blocking either of the two paths as shown in Supplementary Figure 7 and analyzed the results by fitting the parameters.
@@
Supplementary Figure 7: Dependence of $P(H|1)$ and $P(H|2)$ on the $\theta_{GT}$. Graph (a) shows the results measured at the $+$ output port and graph (b) shows the results at the -- output port. The evaluation of the two curves by parameter fitting shows that the condition of $P(H|1)=P(H|2)$ is satisfied for a compensation angle of $0.0407\degree$ for the $+$ output port and an angle of $-0.2739\degree$ for the -- output port.
@@
Graph (a) in Supplementary Figure 7 shows the results measured at the $+$ output port and graph (b) shows the results at the -- output port. The fitting function is
$P(H|1\;\text{or}\;2)=A\cos(n(\theta_{GT}-\theta_{GT0}))+A_{0},$
@@
where the $A$, $n$, $\theta_{GT0}$, $A_{0}$ are fitting parameters. The correct compensation angles should satisfy the condition of $P(H|1)=P(H|2)$ for both output ports. We find that the compensation angle is $0.0407\degree$ for the $+$ output port and $-0.2739\degree$ for the -- output port. The results of $P(H|\pm)$ shown in Fig. 5 were obtained after this adjustment of the direction of the GTs by the corrsponding compensation angles was applied.
@@
Appendix B Background Subtraction
In our experiments, we found that the effective dark counts of our detectors resulted in a constant background of about 400/s for the $+$ port and about 800/s for the -- port. At an intensity of about 110000/s, the Signal-to-Noise ratio (S/N) is 16.5 and 11.7 for the detectors, which correspond to 0.06 and 0.08 in precision, respectively. To achieve a precision of order $10^{-4}$, an intensity of more than $10^{10}$/s would be necessary, but such a high intensity would exceed the operating range of the single photon detectors used in the present experiment. To overcome this problem, we carefully measured the background counts and subtracted them from the counts observed in the experiment reported in this paper.
Supplementary Table 2: Background counts in the measurement time of 100s
$+$port --port
@@ Appendix B Background Subtraction
We performed background runs after each experiment, leaving the configuration unchanged while blocking the source. Separate background runs were performed for both of the experiments where one path was blocked, and for the interference experiment where both paths were open. The measurement time was 100s, performed separately for each polarization at a phase set to $\phi=0$. Supplementary Table 2 shows the summary of the different background counts obtained for the different setups and the different polarization settings. The values in the table were used for background subtraction to obtain the results shown in Figs. 4, 5, and 6.
Appendix C Conditional delocalization
The results presented here are consistent with textbook quantum mechanics and can be predicted theoretically from an analysis of the weak interaction between a system and a quantum probe as shown in [38, 39, 40]. The quantity $A^{2}$ observed when the photon is detected in the output ports of the interferometer associated with constructive ($+$) interference or with destructive (--) interference at a phase shift of $\phi=0$ is determined by the small amount of entanglement between the paths and the polarization generated by the polarization rotation, as explained in [40]. For an input state $\mid\psi\rangle$, the value of $A^{2}$ conditioned by a detection described by a measurement projection $\hat{E}(f)$ is given by
$A^{2}(f)=\frac{\langle\psi\mid\hat{A}\;\hat{E}(f)\;\hat{A}\mid\psi\rangle}{\langle\psi\mid\hat{E}(f)\mid\psi\rangle}$ (4)
@@ Appendix C Conditional delocalization
In the case of an interference measurement, the outcomes $\hat{E}(+)=\mid+\rangle\langle+\mid$ and $\hat{E}(-)=\mid-\rangle\langle-\mid$ are projections of equal superpositions of the states representing the two paths. The commutation relations of $\hat{E}(f)$ and $\hat{A}$ can be given by
$\hat{A}\;\hat{E}(-)\;\hat{A}$ $=$ $\hat{E}(+)$
$\hat{A}\;\hat{E}(+)\;\hat{A}$ $=$ $\hat{E}(-).$ (5)

### v2

@@ 5 Discussion
Our results show how individual photons detected in the outputs of an interferometer are physically delocalized as they propagate through the interferometer, where the concept of delocalization is explained in terms of the empirical evidence provided by local interactions within the paths. By ensuring that the polarization flip rate is the same for all photons localized in only one of the paths, we identify an experimentally observable phenomenon directly associated with the delocalization of the particles between the two paths. As shown in Fig. 5, this phenomenon is consistent with the fact that particles are localized in one of the paths whenever they are detected in that path. However, the situation changes when the particles are detected in the output ports of the interferometer instead. The majority of photons detected in the output port favored by constructive interference experiences a proportional combination of both local effects, indicating a corresponding delocalization of the photons within the interferometer. The stronger the interference effect, the more equal the distribution of each photon between the paths. This effect is consistent with the expectation that interference works best when the intensities of both beams are equal. Photons detected in constructive interference thus behave very much like classical wave intensities. However, their detection in the output port is still a particle-like all-or-nothing event, as demanded by the statistical interpretation of quantum mechanics.
The observed values for the conditional delocalization of a single photon between the paths demonstrate that the relation between statistics and physical quantities cannot be reduced to the naive concept of complementarity suggested in early experiments such as [6]. Instead, there is a causal relation between the quantum fluctuations of $\hat{A}$ and the output statistics of the interferometer. Delocalization causes the photon to exit in the port favored by constructive interference and super-localization causes the rare events of a photon exiting in the other port. Eq.(3) indicates that these fluctuations originate from the same fluctuation of $\hat{A}$ that is also observed in which-path measurements where the squared value of $\hat{A}$ is one in each of the two paths. Fig. 5 shows how the same polarization rotations can indicate either localization or delocalization and super-localization, depending only on the measurement performed after the polarization rotations have been applied. This result illustrates how reality can be measurement dependent. Quantum fluctuations have no intrinsic reality because their specific values depend on their ultimate macroscopic effect. Without such a macroscopic effect in mind, it is not possible to visualize the physics of the system.
This does not mean that the effects of delocalization and super-localization are less real than the localization observed in the direct detection of a particle. In particular, super-localization describes the necessary cause of the unlikely transmission of individual photons to the low probability output that is suppressed by destructive quantum interference. Fig. 6 shows the dependence of super-localization on the phase setting that determines the amount of destructive interference. Maximal values of $A^{2}$ are obtained when destructive interference reduces the output probability to its minimal value. Here, the small polarization rotations used to evaluate super-localization have a noticeable effect, limiting the observable super-localization by introducing a small amount of decoherence. Note that this does not mean that super-localization depends on the interaction used to observe it. In principle, the rotation angle could always be reduced further, demonstrating the inverse relation between super-localization and detection probability for arbitrarily small probabilities.
Super-localization is a fundamental quantum phenomenon associated with the post-selection of destructive quantum interference. We find that this effect is closely related to post-selected metrology, where phase sensitivity can be enhanced beyond the limit set by the quantum uncertainty of the initial state [46, 47, 48]. Indeed, the theory of super-localization in appendix C is very similar to the theory of enhanced phase sensitivity presented in [46, 47]. We are currently working on an experimental realization of a metrology scenario where super-localization is applied to enhance the phase sensitivity of a single photon by a factor corresponding to the enhancement of $A^{2}$ observed in Fig. 6, demonstrating the usefulness of super-localization as a useful resource in quantum metrology.
@@
To summarize our results, we have shown that the physical delocalization of individual particles inside an interferometer can be determined experimentally by using sufficiently weak interactions. The results show that particles are only localized when the final measurement is not sensitive to interference effects. These findings have serious implications for the way in which we think about quantum superpositions. In the absence of empirical evidence to the contrary, it is indeed tempting to claim that each component of a quantum state represents the outcome of a measurement, whether the measurement is performed or not. It is therefore important to recognize that our results demonstrate the fallacy of this counterfactual assumption. In the formalism of quantum mechanics, the role of state vectors and their superpositions is ambiguous. Only an actual experiment can identify the physical meaning of such concepts. Misunderstandings may have arisen from a rather hasty identification of physical properties with the eigenstates of operators that represent them in the formalism. As our previous work has shown[43, 44, 45], the experiment presented here is consistent with the standard formalism of quantum mechanics, and the definition of delocalization used here can be formalized in terms of the appropriate operator ordering (see details in Appendix C). In the present work, we have highlighted the experimental realization because a purely mathematical definition of fundamental concepts would seem arbitrary and carry very little practical meaning. The problem of the conventional approach to quantum mechanics is that its abstractions hide the wide range of possibilities by which we can explore nature. The present work solves this problem for the case of individual particles in an interferometer, paving the way towards a better practical understanding of all quantum phenomena.
@@
Acknowledgements H.F.H. was supported by ERATO, Japan Science and Technology Agency (JPMJER2402); R. F. was supported by JST SPRING, Grant Number JPMJSP2132;
Appendix A. Reduction of systematic effects
A1.1 Elliptical polarization components
@@ Appendix A. Reduction of systematic effects
We found that imperfections of the BS used in the experiment introduced a small amount of elliptical polarization, corresponding to an additional phase shift $\phi_{HV}$ between the H and the V components. If this phase shift is the same for both paths, it causes a systematic error in the $P(H|\pm)$ measurements. To cancel out the elliptical components in the transmitted and reflected output paths, two HWPs placed in the interferometer were individually tilted towards the optical axis. The magnitude of the elliptical component $\langle S_{RL}\rangle$ was directly measured at the ++ and the - ports for path 1 and path 2 by blocking the respective other path in the interferometer. The adjustment of each HWP and the measurement of the elliptical component were iterated to reduce the sum value of the two elliptical polarization components. The measurement results for $\langle S_{RL}\rangle$ obtained with the configuration used in the experiment are summarized in Supplementary Table A1.
Supplementary Table A1: Summary of measurement results
@@
++ port $0759\pm 0.0007 0.0$$604\pm 0.0006 0.0$$0155\pm 0.0009$ $3.52\pm 0.2\degree$
- port $0695\pm 0.0007 0.0$$698\pm 0.0008 0.0$$003\pm 0.0011 0.0$$07\pm 0.2\degree$
@@
The remaining sum value of the elliptical components $\Delta\langle S_{RL}\rangle$ of the transmitted and the reflected paths results in a phase offset in the horizontal axis of Fig. 5. The phase offset $\Delta\phi$ in Supplementary Table A1 was evaluated using the relation $\Delta\langle S_{RL}\rangle/2\theta_{0}$, where the elliptical component can be expressed as $\sin 2\theta_{0}\sin\phi_{HV}\approx 2\theta_{0}\phi_{HV}$. The overall effect of the systematic error is sufficiently small, but it is noticeable that the outcomes of the P(H|)P(H|-) measurement have a lower systematic error than the outcomes of the $P(H|+)$ measurement.
@@
The difference between the systematic errors is particularly noticeable in Fig. 5, where the phase offset $\Delta\phi$ shifts the crossing points of the single-path measurements and the interference measurements. For P(H|)P(H|-), the crossing point is almost precisely at $\phi=90\degree$, whereas for $P(H|+)$ it is not. The mismatch for $P(H|+)$ shown in Fig. 5 (b) is consistent with the offset angle of $\Delta\phi=3.52\degree$, which roughly corresponds to 2/3 of one step of the data points shown in Fig. 5 (b).
A1.2 Additional polarization rotation by the final BS
@@
Another systematic effect originates from an additional polarization rotation caused by an unintended polarization dependence of transmission and reflection at the BS. This additional polarization rotation introduces a difference in the magnitude of the rotation observed in the two output ports that is not related to the delocalization of the photon. To avoid this systematic effect, the orientations of the GTs in the output ports were slightly adjusted so that the values of $P(H|1)$ and $P(H|2)$ observed when the other path was blocked had the same values in both output ports. This adjustment is necessary to satisfy the condition that the magnitude of the polarization rotation is determined by the opposite polarization rotations in the paths only, with no systematic dependence on the output path. To evaluate the compensation angle, we measured the dependence of $P(H|1)$ and $P(H|2)$ on the angle of the GT $\theta_{GT}$ by blocking either of the two paths as shown in Supplementary Figure A1 and analyzed the results by fitting the parameters.
@@
Supplementary Figure A1: Dependence of $P(H|1)$ and $P(H|2)$ on the $\theta_{GT}$. Graph (a) shows the results measured at the ++ output port and graph (b) shows the results at the - output port. The evaluation of the two curves by parameter fitting shows that the condition of $P(H|1)=P(H|2)$ is satisfied for a compensation angle of $0.0407\degree$ for the ++ output port and an angle of $0.2739\degree$ for the - output port.
@@
Graph (a) in Supplementary Figure A1 shows the results measured at the ++ output port and graph (b) shows the results at the - output port. The fitting function is
$P(H|1\;\text{or}\;2)=A\cos(n(\theta_{GT}-\theta_{GT0}))+A_{0},$ (A1)
@@
where the $A$, $n$, $\theta_{GT0}$, $A_{0}$ are fitting parameters. The correct compensation angles should satisfy the condition of $P(H|1)=P(H|2)$ for both output ports. We find that the compensation angle is $0.0407\degree$ for the ++ output port and $0.2739\degree$ for the - output port. The results of $P(H|\pm)$ shown in Fig. 5 were obtained after this adjustment of the direction of the GTs by the corrsponding compensation angles was applied.
@@
Appendix B. Background Subtraction
In our experiments, we found that the effective dark counts of our detectors resulted in a constant background of about 400/s for the ++ port and about 800/s for the - port. At an intensity of about 110000/s, the Signal-to-Noise ratio (S/N) is 16.5 and 11.7 for the detectors, which correspond to 0.06 and 0.08 in precision, respectively. To achieve a precision of order $10^{-4}$, an intensity of more than $10^{10}$/s would be necessary, but such a high intensity would exceed the operating range of the single photon detectors used in the present experiment. To overcome this problem, we carefully measured the background counts and subtracted them from the counts observed in the experiment reported in this paper.
Supplementary Table A2: Background counts in the measurement time of 100s
++port -port
@@ Appendix B. Background Subtraction
We performed background runs after each experiment, leaving the configuration unchanged while blocking the source. Separate background runs were performed for both of the experiments where one path was blocked, and for the interference experiment where both paths were open. The measurement time was 100s, performed separately for each polarization at a phase set to $\phi=0$. Supplementary Table A2 shows the summary of the different background counts obtained for the different setups and the different polarization settings. The values in the table were used for background subtraction to obtain the results shown in Figs. 4, 5, and 6.
Appendix C. Conditional delocalization
The results presented here are consistent with textbook quantum mechanics and can be predicted theoretically from an analysis of the weak interaction between a system and a quantum probe as shown in [43, 44, 45]. The quantity $A^{2}$ observed when the photon is detected in the output ports of the interferometer associated with constructive (++) interference or with destructive (-) interference at a phase shift of $\phi=0$ is determined by the small amount of entanglement between the paths and the polarization generated by the polarization rotation, as explained in [45]. For an input state $\mid\psi\rangle$, the value of $A^{2}$ conditioned by a detection described by a measurement projection $\hat{E}(f)$ is given by
$A^{2}(f)=\frac{\langle\psi\mid\hat{A}\;\hat{E}(f)\;\hat{A}\mid\psi\rangle}{\langle\psi\mid\hat{E}(f)\mid\psi\rangle}$ (A2)
@@ Appendix C. Conditional delocalization
[…]